Charlie has scores of 88, 95, and 89 on the first three tests. What score must he make on the fourth test so that the average of all four tests is exactly 93?
a.91.25 b.93 c.95 1/3 d.97 e.100


Found 2 solutions by AAfter Search, mananth:
Answer by AAfter Search(61)

Let Charlie scores N in the fourth test to get an average of 93 marks.
Hence, (88 + 95 + 89 + N)/4 = 93
=> 272 + N = 4 x 93 = 372
=> N = 372 - 272 = 100
Hence, Charlie must score 100 in the fourth test to get an average of 93 marks in all four tests.

Answer by mananth(16946)

charlie's first 3 scores = 88,89,95
Total of these scores = 272
..
He needs an average of 93 in 4 tests
so total of 4 tests will be 93 *4=372
..
The difference of the totals will give the score for the 4th. test to achieve 93 average
372-272=100